bobthebuilder696968:

Given that f(x) = x2 − 4x − 3 and g(x) = the quantity of x plus three, over four , solve for f(g(x)) when x = 9. (1 point) −6 3 6 9

dude:

\(g(x)=\frac{x+3}{4}\)

dude:

\(f(g(x))\) Basically means to substitute g(x) into f(x)

dude:

\(f(g(x))=\boxed{\frac{x+3}{4}}^2 − 4(\boxed{\frac{x+3}{4}}) − 3\)

bobthebuilder696968:

@dude ahh okay thank you